OpenStudy (anonymous):

The slope of a line is -1/3, and the y-intercept is 10/3. What is the equation of the line written in general form? 10x + 3y - 1 = 0 x + 3y + 10 = 0 x + 3y - 10 = 0

OpenStudy (midhun.madhu1987):

First convert it into the slope-intercept form: y = mx + b where m= slope and y-intercept what u get?

OpenStudy (anonymous):

I got -1/3 + 10/3 but that doesn't make any sense.

OpenStudy (midhun.madhu1987):

y = mx + b plug in -1/3 for m and 10/3 for b

OpenStudy (anonymous):

y = -1/3x + 10/3.

OpenStudy (midhun.madhu1987):

yeah... now we need to make rearrangements and convert it to the Standard form which is: ax + by + c = 0

OpenStudy (midhun.madhu1987):